OpenStudy (help_people):

If h(x) = x − 1 and j(x) = −4x, solve h[j(5)] and select the correct answer below. −21 −20 −19 −16

OpenStudy (jtvatsim):

First find j(5). Remember that this is just "shorthand" that says "Plug x = 5 into the j equation". We get j(5) = -4*5 = -20. Now, we can find h[j(5)], which means "Plug x = j(5) into the h equation". Well, we know that j(5) is just -20. So we will plug in x = -20 into the h equation. This gives us: h[j(5)] = h[-20] = -20 - 1 = -21. This is our final answer.

OpenStudy (help_people):

wwow ty so much thanks for explaining it! :)